# Why the Agent Made that Decision: Explaining Deep Reinforcement Learning with Vision Masks

### v1

@@ 4.1 Assessment of Explanation Fidelity
[…]
Values of these fidelity metrics are shown in Table 2. VisionMask consistently scores higher than LIME and RISE. Greydanus and SARFA have a higher precision than VisionMask, however, a much lower recall and F1-score. This is probably because they included almost all pixels from the visual input in the importance map. They also have a lower accuracy compared to VisionMask.
Refer to caption Refer to caption
Original state VisionMask
Refer to caption Refer to caption
Insertion curve Deletion curve
Figure 2: Examples of insertion and deletion AUC curve
Env Method 0.25 0.5 1.0
$\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins
Mario VisionMask 0.2044 ±0.1891 0.4156 ±0.2937 0.1813 ±0.1730 0.4142 ±0.2788 0.1714 ±0.1577 0.4216 ±0.2622
LIME 0.3447 ±0.2664 0.2314 ±0.2189 0.3231 ±0.2568 0.2446 ±0.2232 0.2693 ±0.1958 0.3024 ±0.2107
RISE 0.2750 ±0.2510 0.2940 ±0.2624 0.2331 ±0.2326 0.3317 ±0.2761 0.1892 ±0.1897 0.3796 ±0.2678
Greydanus 0.5708 ±0.2109 0.1963 ±0.1658 0.5671 ±0.2146 0.1952 ±0.1650 0.5634 ±0.2169 0.2018 ±0.1691
SARFA 0.5579 ±0.2223 0.2071 ±0.1756 0.5565 ±0.2235 0.2117 ±0.18 0.5556 ±0.2246 0.2143 ±0.1866
Enduro VisionMask 0.3693 ±0.2854 0.3952 ±0.3004 0.3589 ±0.2911 0.3979 ±0.2995 0.3491 ±0.2948 0.4021 ±0.2941
LIME 0.3780 ±0.2886 0.3636 ±0.3025 0.3723 ±0.2914 0.3680 ±0.3002 0.3629 ±0.2937 0.3799 ±0.2883
RISE 0.4081 ±0.2818 0.3585 ±0.3066 0.3970 ±0.2861 0.3714 ±0.2938 0.3759 ±0.2912 0.3906 ±0.2785
Greydanus 0.3762 ±0.2988 0.3312 ±0.3194 0.3626 ±0.3018 0.3317 ±0.3192 0.3457 ±0.3075 0.3329 ±0.3188
SARFA 0.3472 ±0.3060 0.3327 ±0.3188 0.3378 ±0.3114 0.3336 ±0.3186 0.3331 ±0.3157 0.3340 ±0.3186
Seaquest VisionMask 0.0721 ±0.0291 0.3256 ±0.2596 0.0684 ±0.0289 0.3331 ±0.2650 0.0652 ±0.0268 0.3430 ±0.2635
LIME 0.1982 ±0.1871 0.1397 ±0.1662 0.1909 ±0.1865 0.1425 ±0.1672 0.1434 ±0.1214 0.2141 ±0.1812
RISE 0.1198 ±0.0797 0.2422 ±0.2560 0.0987 ±0.0658 0.2638 ±0.2654 0.0809 ±0.0505 0.2896 ±0.2660
Greydanus 0.2939 ±0.2381 0.0658 ±0.0345 0.2165 ±0.1834 0.0708 ±0.0542 0.1439 ±0.0980 0.0788 ±0.0825
SARFA 0.1133 ±0.0732 0.0819 ±0.0978 0.0888 ±0.0418 0.0851 ±0.1104 0.0764 ±0.0307 0.0869 ±0.1177
MsPacman VisionMask 0.2251 ±0.1051 0.6350 ±0.3056 0.2037 ±0.1255 0.6231 ±0.3337 0.1956 ±0.1217 0.6278 ±0.2778
LIME 0.4831 ±0.2700 0.1901 ±0.1708 0.4348 ±0.2455 0.2346 ±0.1875 0.3265 ±0.1801 0.3771 ±0.1890
RISE 0.4270 ±0.3140 0.2299 ±0.1220 0.3512 ±0.2995 0.2937 ±0.1325 0.2621 ±0.2273 0.4577 ±0.1394
Greydanus 0.5669 ±0.2476 0.1285 ±0.0646 0.4903 ±0.2518 0.1336 ±0.0843 0.3905 ±0.2281 0.1417 ±0.1029
SARFA 0.3511 ±0.2528 0.1463 ±0.1168 0.2524 ±0.1573 0.1531 ±0.1328 0.1880 ±0.0873 0.1561 ±0.1425
Table 3: Insertion and deletion metrics for fractions 0.25, 0.5, and 1.0. The symbol $\uparrow$ (resp. $\downarrow$) indicates that higher (resp. lower) values are better. We report the mean $\pm$ standard deviation over testing samples.
4.2 Assessment of Effectiveness
An effective explainer should precisely identify the exact input regions that contribute to the expert decision, without including irrelevant areas. To quantify the effectiveness of explanations, We adopt the insertion and deletion metrics from RISE (Petsiuk, Das, and Saenko 2018). The rationale behind these metrics is as follows: If certain pixels in the visual state are crucial to the action (i.e., the true cause of the action), deleting them from the state or inserting them into the background should result in a corresponding decrease or increase in the probability of the specific action. By plotting the relationship between the probability of the target action and the number of pixels added or removed, we obtain an insertion curve and a deletion curve. The Area Under the Curve (AUC) of those plots represents the effectiveness of the explanation. A higher AUC for the insertion curve and a lower AUC for the deletion curve stand for more effective explanation. The order of insertion and deletion affects the result. We sort the pixels based on their values in the importance map and insert or delete them in descending order of their importance. Figure 2 shows an example of insertion and deletion curves with corresponding AUC values displayed at the top right corner. See Appendix D for the pseudocode of each metric.
It is worth noting that the effectiveness varies as different number of pixels are inserted or deleted from the original input. Instead of showing the actual curve for each explainer under different games, we report the AUC under 25%, 50%, and 100% level of insertion and deletion. The results are given in Table 3. VisionMask outperforms all other methods across all environments, with an average insertion score that is 14.9% higher and a deletion score that is 8.8% lower compared to the other methods.
4.3 Counterfactual Explanations
Original state VisionMask Counterfactual
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Move left Jump right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Left Left + acceleration
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Down Fire Down right
Figure 3: Counterfactual analysis with VisionMask.
By removing the most important regions identified by VisionMask from the visual input, we can create scenarios for counterfacutal analysis. Table 3 provides three examples of this approach. In the first example, VisionMask correctly identified the Piranha Plants as the reason the agent chose to move left. After removing the Piranha Plant by replacing the highlighted pixels with reference value and run the modified image by the agent again, we observed that the agent’s action changed to jumping right, as the threat from the Piranha Plant no longer existed. In the second example, VisionMask correctly identified the edge of the road and a car in front as the reason why the agent chose to move left (with current speed). After we removed the car in front the agent chose to move left with acceleration. In the third example, VisionMask identified the shark in the bottom left as the cause of the downward and fire action. When the corresponding shark was removed, the agentno longer chose to fire. Appendix I, J and K contain more examples. We believe such explanations can further help users to understand the RL agent’s behavior.
@@ 5 Conclusion
We presented VisionMask, an agent-agnostic DRL explanation model trained in self-supervised contrastive learning. VisionMask generates explanations with higher fidelity and better effectiveness compared to existing attribute-based methods. It is our future plan to extend this approach to multi-modality input and couple the visual explanation generated by the VisionMake with other information such as agent’s long-term goals and future rewards.
Acknowledgements
This research is partially supported by the Air Force Office of Scientific Research (AFOSR), under contract FA9550-24-1-0078. The paper was received and approved for public release by AFRL on May 28th 2024, case number AFRL-2024-2908 . Any Opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of AFRLor its contractors.
Appendix A Technical Appendix
Appendix B Meanings of actions across four environments
In Table 4, we present the list of actions and their corresponding meanings for the four environments evaluated in the experiment section.
Value Meaning Value Meaning
@@ 8 Accelerate and move left
(a) Mario environment (b) Atari Enduro environment
Value Meaning Value Meaning
@@ 17 Fire and move down-left
(c) Atari Seaquest environment (d) Atari MsPacman environment
Table 4: Action meanings for SMB (Nguyen 2020), Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016)
Appendix C Reference values used across different environments
In our experiment, we found that the PPO agents we adopted from Stable Baselines3 (Raffin et al. 2021) and Mario-PPO (Nguyen 2020) are sensitive to reference values. We tried using a constant value (zero) and blur as reference values but did not obtain clear and reasonable importance masks. However, using reference values that are visually similar to the environment background gave us better results. In Figure 4, we list the reference values we used in each environment. Specifically, for the SMB and Atari Enduro environments, we use the background color of the content in each state as the reference value. Note that the reference value for different states may vary, as the background content of each state can be different. For the Atari Seaquest environment, since all the states share the same background, we use a single reference value for each state, which combines the empty sea background and the oxygen bar background. For the Ms. Pacman environment, we simply used a constant value of the background color blue as the reference value.
Figure 4: Reference values for four environment evaluated. The content column represents a sub-figure of the original state that contains the objects and background. The last column displays the corresponding reference value for each environment.
Env Original state Content Reference value
SMB [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix D Algorithms for insertion and deletion metrics
Algorithms 1-2 present the pseudocode for evaluating the insertion and deletion metrics of (Petsiuk, Das, and Saenko 2018).
@@ Appendix D Algorithms for insertion and deletion metrics
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels inserted $\alpha$, Number of pixels $N$
@@
3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
@@
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels deleted $\alpha$, Number of pixels $N$
@@
3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
@@
Appendix E Object importance percentage calculation
[…]
Appendix F Time efficiency
To assess the time efficiency of mask generation, we compare the total time taken for mask generation using VisionMask against that of two baseline methods on testing datasets from four environments. Note that the time we measure does not include the training time for VisionMask. We can see that LIME and RISE are slower than VisionMask. All the experiments are conducted on A40 GPU.
Env VisionMask LIME RISE
SMB 00:00:10 25:45:02 00:51:41
Enduro 00:00:10 1:05:19 00:08:14
Seaquest 00:00:12 1:34:49 00:10:58
MsPacman 00:00:10 3:02:28 00:08:59
Table 5: Time taken to generate importance masks for three methods across four environments. The time format is hh:mm:ss. The testing dataset for SMB contains 970 states, while Atari Enduro, Seaquest, and Ms. Pacman contain 4000 states each.
Appendix G Experimental details
Dataset details
Recall from Section 3.1 that we collected the expert dataset $\pi_{E}$ with the agent being explained as expert in four envrionments. In the SMB environment(Kauten 2018), $\pi_{E}^{\text{Mario}}$ comprises a total of 9,692 state-action pairs. In Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016) environment, $\pi_{E}^{\text{Atari}}$ includes 10,000 total state-action pairs. We utilize a 80%-10%-10% train-valid-test split on all expert dataset $\pi_{E}$.
Pre-processing data
We transform the RGB input to greyscale with the following function then resize it to $84\times 84$:
1def rgb_to_grayscale(img, num_output_channels: int = 1):
2 if img.ndim < 3:
3 raise TypeError(f"Input image tensor should have at least 3 dimensions, but found {img.ndim}")
4 # _assert_channels(img, [1, 3])
5
6 if num_output_channels not in (1, 3):
7 raise ValueError("num_output_channels should be either 1 or 3")
8
9 if img.shape[-3] == 3:
10 r, g, b = img.unbind(dim=-3)
11 # This implementation closely follows the TF one:
12 # https://github.com/tensorflow/tensorflow/blob/v2.3.0/tensorflow/python/ops/image_ops_impl.py#L2105-L2138
13 l_img = (0.299 * r + 0.587 * g + 0.114 * b).to(img.dtype)
14 l_img = l_img.unsqueeze(dim=-3)
15 else:
16 l_img = img.clone()
17
18 if num_output_channels == 3:
19 return l_img.expand(img.shape)
20
21 return l_img
Model details
We utilize the segmentation model DeeplabV3 as the $f_{\theta}$ model in Section 3. See the Github repo of this model from  [^1] https://github.com/pytorch/pytorch. As mentioned in the Experiment section 4, we adopt the pretrained PPO model from Stable baselines 3 (Raffin et al. 2021) as the agent to be explained. Check the Hugging Face repositories for detailed information on the PPO agents for Enduro  [^2] https://huggingface.co/sb3/ppo-EnduroNoFrameskip-v4, Seaquest  [^3] https://huggingface.co/sb3/ppo-SeaquestNoFrameskip-v4, and MsPacman  [^4] https://huggingface.co/sb3/ppo-MsPacmanNoFrameskip-v4.
Training details
We conduct the training process on a server with an A40 GPU, using 50 epochs for SMB, 36 epochs for Enduro, 52 epochs for Seaquest, and 136 epochs for Ms. Pac-Man. The optimization is done using Adam. See Table 6 for hyperparameters.
Hyperparameter value
Learning rate 0.00001
Training batch size 16
Testing batch size 1
Validating batch size 16
$\lambda_{avg}$ 0.3
$\lambda_{e}$ 1.0
$\lambda_{\text{smooth}}$ 1.0
$\lambda_{L2}$ 0.01
Table 6: Hyperparameters
Three random seeds used for training: 42, 13, 62.
Appendix H Code
Code can be found at: https://anonymous.4open.science/r/I6eV3cksE5AgZ7-EB88
Appendix I Counterfactual explanations of Super Mario Bros environment
In this section, we show more counterfactual explanations on the Super Mario Bros(Kauten 2018) environment. In Table 7 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detail the counterfactual explanations in Table 7 below:
Note that all the object importance percentages are calculated following E.

### v2

@@ 4.2 Quantitative Analysis
Insertion and deletion measure the impact on the target action probability by progressively adding or removing pixels from the original input based on the descending order of their attribution scores, with the most important pixels being inserted/deleted first. The inserted/deleted input image will be processed by the expert agent again to obtain the probability of the target action. By plotting the probability against the number of pixels added or removed, we obtain an insertion and a deletion curve. The insertion/deletion performance is measured by the area under the curve (AUC) of the insertion/deletion curves. A larger (smaller) AUC for the insertion (deletion) curve means that including (removing) the important pixels identified by the explainer can effectively increase (reduce) the probability of the target action. Hence a larger (smaller) AUC of insertion (deletion) curve indicates more accurate attribution prediction. The pseudo-code of the detailed information of deletion and insertion could be found in Appedix E.
For robustness, we report the Local Lipschitz Estimate (LLE) scores Alvarez Melis and Jaakkola (2018), which quantify the local smoothness of explanations by estimating the Lipschitz constant within a specific neighborhood. The Lipschitz constant measures the maximum rate of change of the function, ensuring explanation does not vary too rapidly within the state’s neighborhood. Given state $s_{i}$ and neighborhood size $\epsilon$, LLE defined as
$\hat{L}(s_{i})=\underset{s_{j}\in\mathcal{N}_{\epsilon}(s_{i})}{\text{argmax}}\frac{\|f_{\theta}(s_{i})-f_{\theta}(s_{j})\|_{2}}{\|s_{i}-s_{j}\|_{2}}$
where $\mathcal{N}_{\epsilon}(s)=\{s^{\prime}\in X\mid\|s-s^{\prime}\|\leq\epsilon\}$.
We evaluate complexity with Sparseness Chalasani et al. (2020), which uses the Gini index on the vector of absolute attribution values sorted in non-descending order. Sparseness ensures that features genuinely influencing the output have substantial contributions, while insignificant or only slightly relevant features should have minimal contributions. A higher Sparseness indicates more contrastive attribution values and hence more understandable explanations  Chalasani et al. (2020).
Results. In Table 1, our VisionMask achieves the best performance in terms of faithfulness (i.e., Acc, Del, Ins) in all testing environment except SMB, where its deletion score is slightly lower than RISE. However, RISE has significantly lower insertion score and accuracy in this environment. Hence, our explanations are more aligned with the agent’s decision-making process. Moreover, compared to, Explainer Stalder et al. (2022), VisionMask exhibits much higher faithfulness, which suggests the effectiveness of action contrastive learning. Overall, learning-based model performs better compared to perturbation based approach.
In terms of robustness, RISE achieves the best performance in terms of LLE score across all settings. This is because, as a perturbation-based method, RISE uses the same type of perturbation that we used to generate the neighbor image for LLE score calculation. Hence the perturbation almost has no impact to it. On the other hand, RISE has the lowest sparsity, which means the attribution predicted by RISE are evenly distributed and contains little information. For Complexity, benefiting from the binary mask of LIME which is much more sparse compared to other baselines, LIME has the highest sparseness score. However, it has the worst performance in terms of LLE score. Overall, VisionMask achieves the best balance between the robustness and sparseness. See Appendix for the Radar map.
Metrics Stalder L1
Acc. 95.6 95.9
Ins. $\uparrow$ 64.3 67.6
Del. $\downarrow$ 21.1 20.4
LLE $\downarrow$ 56.6 62.6
Sp. $\uparrow$ 73.8 62.3
Hours 0.72 0.68
Speedup - $1.1\times$
Sp./LLE 1.3 1.0
(a) Area size regularization. L1 is better for faithfulness and faster. Time is estimated by training 10 epochs.
Metrics FCN UNet DeepLabv3
Acc. 94.2 94.7 96.0
Ins. $\uparrow$ 56.5 56.8 67.5
Del. $\downarrow$ 19.6 19.7 20.4
LLE $\downarrow$ 100.0 86.3 62.6
Sp. $\uparrow$ 58.0 19.8 62.3
Sp./LLE 0.58 0.23 0.99
(b) $f_{\theta}$ model. DeepLabv3 achieves better results.
Metrics $0.1$ $0.2$ $0.4$
Acc. 95.9 94.5 95.8
Ins. $\uparrow$ 57.0 56.9 57.0
Del. $\downarrow$ 19.5 19.5 19.5
LLE $\downarrow$ 96.0 93.8 99.5
Sp. $\uparrow$ 79.0 58.6 48.1
Sp./LLE 0.82 0.62 0.48
(c) Max area size $a_{\text{max}}$. Maximum area size of $0.1$ produces favorable results.
Metrics background black mean blur
Acc. 97.8 95.9 97.6 94.4
Ins. $\uparrow$ 67.5 51.7 54.5 56.9
Del. $\downarrow$ 20.4 16.8 18.2 25.0
LLE $\downarrow$ 62.6 54.4 92.6 99.6
Sp. $\uparrow$ 62.3 28.5 45.5 56.6
Sp./LLE 1.0 0.52 0.49 0.57
(d) Reference value. Using background produce better faithfulness values.
Comp. C. +NE +NE+L1 +NE+L1+TV
Acc. 96.7 96.5 95.9 95.6
Ins. $\uparrow$ 56.7 56.5 67.6 64.6
Del. $\downarrow$ 23.7 23.8 20.4 20.7
LLE $\downarrow$ 99.3 96.0 62.6 68.8
Sp. $\uparrow$ 24.0 23.8 62.3 60.5
Sp./LLE 0.24 0.25 0.99 0.88
(e) Component. Negative Entropy (NE) with L1 area regularization performs better in terms of faithufulness. Contrastive-only (C.), Stalder Area Loss (S.), and Total Variation (TV.)
Table 2: VisionMask ablation experiments on SMB. Default settings are marked in gray.
4.3 Qualitative Analysis
In this section, we conduct two qualitative analyses across two settings, visual comparision and human studies.
Visual Comparision. We present example explanations from three environments, SMB, Enduro, and Seaquest, in Figure 2, along with some counterfactual analysis generated from the explanations provided by VisionMask. The examples show that both LIME and RISE fail to generate interpretable explanations. LIME’s superpixels are too large to capture the specific regions, while RISE’s explanations include almost all pixels. Explainer generates more accurate and interpretable explanations compared to LIME and RISE. This suggests that purely perturbation-based approaches may fail in RL due to the high dynamics of environments and the lack of learning ability. Greydanus et al. (2018) and SARFA generate better explanations than LIME and RISE but often focus on irrelavent objects or background. In contrast, VisionMask accurately highlights the relevant regions, providing sharp explanations that are both accurate and interpretable. By removing some of the regions highlighted by the VisionMask, counterfactual analysis could be performed to answer questions like: ”Why this action instead of that one?”
Figure 3: Human studies. Comparison of total responses and positive responses across questions.
Human Studies. To assess whether the saliency maps improve humans’ understanding of the agent’s decisions, we conduct human studies for SMB, Enduro and Seaquest. We present 63 participants with 9 state-action pairs accompanied by the saliency maps generated by VisionMask. Participants are then asked whether the saliency maps help them better understand the decision-making process of the RL agents. We record the total number of responses as well as the number of positive responses, as shown in Figure 3. Over 89% of participants find the saliency maps generated by VisionMask helpful in understanding the agents’ decisions. Details about this study can be found in the Appendix D.
4.4 Ablation Studies
Table 2 shows the results for ablation studies. In all tables, the column marked in gray represents the default setting of VisionMask used to generate results in previous sections.
Area size regularization. In this experiment, we replace the default L1 area regularization with the Min-Max Area Loss from Explainer Stalder et al. (2022). Without this constraint, Explainer tends to produce masks where all values are one in order to achieve the best performance. Explainer first vectorizes and sorts the mask elements, and then applies penalties to pixels falling outside the range $[a,b]$, where the hyperparameters $a$ and $b$ specify the minimum and maximum allowable area sizes. Instead, VisionMask uses L1 area regularization to constrain the maximal area, as we do not wish to impose constraints on the minimum area size. To implement action contrastive learning, we need to generate masks for actions that are not selected, and these masks should be allowed to be all zeros if necessary. As shown in Table 2a, using L1 area regularization improves the faithfulness of the explanation and increases training speed.
Max area size. We further vary the maximum permitted area size $a_{\text{max}}$ to evaluate its impact. As shown in Table 2c, reducing $a_{\text{max}}$ significantly increases the sparseness as it constrains the model to focus on the most discriminative regions without compromising other performance metrics. The $0.1$ setting achieves the best results, as expected.
$\mathbf{f_{\theta}}$ model. In Table 2b, we compare the performance of three different versions of VisionMask with three different segmentation models: our default DeepLabv3 Chen et al. (2017), Fully Convolutional Network (FCN)Long et al. (2015), and UNetRonneberger et al. (2015). The results show that DeepLabv3 achieves a better performance.
Reference Value. We evaluate the impact of hyperparameter $r$ in Equation 1 by setting the reference value to background, black, mean, and blur. The ”black” reference sets $r$ to 0, the ”mean” reference sets $r$ to the RGB mean value of the entire dataset, and the ”blur” reference applies Gaussian blur (kernel is 39, $\sigma=15$) to the image and use the blurred pixel value as $r$. Table 2d compares the performance of VisionMask with these four different reference values. We choose the background as reference value as it provides a more balanced performance across all metrics. Details about the background reference value are provided in Appendix F.
Loss Function Components. In Table 2e, we evaluate the impact of different components in the loss function on the performance of VisionMask. The results show that, although adding negative entropy and L1 area regularization slightly reduces accuracy, it improves the insertion and deletion scores and significantly boosts the Sp./LLE ratio. In contrast, adding total variation to produce smoother masks has minimal impact on overall performance.
@@ 5 Conclusion
We presented VisionMask, an agent-agnostic DRL explanation model trained in self-supervised contrastive learning. VisionMask generates explanations with higher fidelity and better effectiveness compared to existing attribute-based methods. It is our future plan to extend this approach to multi-modality input and couple the visual explanation generated by the VisionMask with other information such as agent’s long-term goals and future rewards.
Acknowledgments
This research is partially supported by the Air Force Office of Scientific Research (AFOSR), under contract FA9550-24-1-0078, and NSF award CNS-2148253. The paper was received and approved for public release by Air Force Research Laboratory (AFRL) on May 28th 2024, case number AFRL-2024-2908 . Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of AFRL or its contractors.
Appendix A Agent details
config value
learning rate 0.001
batch size 32
$\epsilon$ 0.5
$\gamma$ 0.99
buffer size 50000
target update interval 1000
Table 3: VizDoom DQN
config value
learning rate 0.0005
batch size 32
$\epsilon$ 0.5
$\gamma$ 0.8
buffer size 15000
target update interval 50
Table 4: Highway-env DQN
As mentioned in the section 4, we adopt the pretrained PPO models from stable-baselines3 Raffin et al. [2021] for Enduro, Seaquest and MsPacman. Check the Hugging Face repositories for detailed information on the PPO agents for Enduro [^1] https://huggingface.co/sb3/ppo-EnduroNoFrameskip-v4, Seaquest [^2] https://huggingface.co/sb3/ppo-SeaquestNoFrameskip-v4, and MsPacman [^3] https://huggingface.co/sb3/ppo-MsPacmanNoFrameskip-v4. For SMB, we use SMB-PPONguyen [2020] [^4] https://github.com/vietnh1009/Super-mario-bros-PPO-pytorch. Inputs for above agents are 4 continuous RGB frames without frame skipping of shape $84\times 84$. For VizDoom, we train the DQN agent from scratch. The agent consists of four convolutional layers, each with a kernel size of 3, stride of 1, and filter sizes ranging from 64 to 512 (64, 128, 256, 512). Inputs are 5 greyscale frames wihout frame skipping of shape $60\times 40$. The training setting is in Table 3. For Highway-env, we train the DQN agent of 2 convolutional layers (kernel size 3, stride 1 and filter size 32). Inputs are 4 greyscale frames without frame skipping of shape $128\times 64$. Details of training is in Table 4.
Appendix B Datasets
env train val test state dim state format actions
SMB 7754 970 969 $4\times 256\times 240$ RGB 7
Enduro 8000 1000 1000 $4\times 160\times 210$ RGB 9
Seaquest 8000 1000 1000 $4\times 160\times 210$ RGB 9
MsPacman 8000 1000 1000 $4\times 160\times 210$ RGB 9
VizDoom 13336 1667 1667 $5\times 60\times 40$ Grey 7
Highway-env 15940 1992 1993 $4\times 64\times 128$ Grey 5
Table 5: Datasets details of six environments
With the aforementioned agents, we collect total state-action pairs of 9693 for SMB in 31 different worlds and stages following Nguyen [2020], 10000 for Enduro, Seaquest and MsPacman, 16670 for VizDoom and 19925 for Highway-env. See Table 5 for details. The datasets is publicly available  [^5] https://huggingface.co/datasets/0pow0/visionmask_datasets
Mario Meaning Enduro Meaning
@@ 8 Accelerate and move left
Seaquest Meaning MsPacman Meaning
@@ 17 Fire and move down-left
VizDoom Meaning Highway-env Meaning
0 TURN LEFT 0 LANE LEFT
1 TURN RIGHT 1 IDLE
2 ATTACK 2 LANE RIGHT
3 MOVE LEFT 3 FASTER
4 MOVE RIGHT 4 SLOWER
5 MOVE FORWARD
6 MOVE BACKWARD
Table 6: Action meanings for SMB, Enduro, Seaquest and MsPacman
Appendix C Actions
In Table 6, we present the list of actions and their corresponding meanings for the six environments evaluated in the experiment section.
Appendix D Example of the human studies
Figure 4: Example interface for our human studies
In Figure 4, we present a screenshot of the interface used for our human studies. We selected three examples from SMB, Enduro, and Seaquest. Participants were asked to provide Yes/No responses based on whether the saliency map’s explanation aligned with their understanding and if it was reasonable.
Appendix E Algorithms for insertion and deletion metrics
Algorithms 1-2 present the pseudocode for evaluating the insertion and deletion metrics of Petsiuk et al. [2018].
@@ Appendix E Algorithms for insertion and deletion metrics
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels inserted $\alpha$, Number of pixels NN
@@
[…]
@@
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels deleted $\alpha$, Number of pixels NN
@@
[…]
@@
Figure 5: Reference values for four environment. The content column represents a sub-figure of the original state that contains the objects and background. The last column displays the corresponding reference value for each environment.
Env Original state Content Reference value
SMB [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix F Background reference values used across different environments
In Figure 5, we list the reference values we used in each environment. Specifically, for SMB and Enduro, we use the background color of the content in each state as the reference value. For Seaquest environment, since all the states share the same background, we use a single reference value for each state, which combines the empty sea background and the oxygen bar background. For MsPacman, we simply used a constant value of the background color blue as the reference value. For VizDoom, we use black background; And we use the default grey background for Highway-env
Appendix G Radar map for quantitative results
Figure 6: Radar map of VisionMask and five baselines in the six environments.
We present the radar map in Figure 6 to better visualize the quantitative results. Each metric was averaged across all environments. For deletion (Del.) and LLE, we calculated the complement by subtracting the value from 100%. VisionMask demonstrates better faithfulness (shown in red) compared to other baselines.
Figure 7: Additional qualitative examples compared with five baselines in three environments. (a) Human explanation: ”Mario jumps to the right to get the Koopa on the ground.” VisionMask correctly identified the Koopa. Counterfactual analysis shows that removing the two Koopas on the ground changes the action from ’Jump right’ to ’Right.’ Note that ’Right’ cannot reach the Koopa. (b) Human explanation: ”Mario jumps to the right to land on the platform.” VisionMask correctly identified the platform. Counterfactual analysis shows that removing the platform changes the action from ’Jump right’ to ’Jump.’ (c) Human explanation: ”The agent moves to the right while maintaining its current speed because it is blocked by the car in front.” VisionMask correctly identified the agent’s current location, the front car, and the track. Counterfactual analysis shows that removing the front cars allows the agent to accelerate instead of choosing ’Right + accelerate.’ (d) Human explanation: ”The agent moves to the right to avoid a crash with the car on the left.” VisionMask correctly identified the agent’s current location and the two cars on the left. Counterfactual analysis shows that removing the cars on the left enables the agent to choose ’Left + accelerate’ instead of ’Right + accelerate.’ (e) Human explanation: ”The agent moves up to rescue the diver.” VisionMask correctly identified the relative locations of the agent and the diver, as well as the oxygen bar. Counterfactual analysis shows that removing the diver causes the submarine to stop moving up. (f) Human explanation: ”The agent stays on the surface to wait for the oxygen to fully recharge.” VisionMask correctly identified the oxygen bar. Counterfactual analysis shows that removing the rightmost oxygen bar causes the submarine to start operating.
Appendix H Additional qualitative explames
We present further qualitative examples comparing VisionMask to baselines in SMB, Enduro and Seaquest environments in Figure 7.
Appendix I Object importance percentage calculation
Attention percentages paid to objects are calculated by summarizing the attention map on target objects. Existing object detection models, trained on real-world images, perform unsatisfactorily in gaming environments. Furthermore, most object detection models use bounding boxes to detect objects, which are not suitable for fine-grained attention percentage calculations. To address these challenges, we use a segmentation model as an alternative approach. Specifically, we employ Segment Anything (SAM) Kirillov et al. [2023], which automatically generates masks indicating potential target objects, denoted as $\mathbf{m}^{\text{SAM}}\in{\{0,1\}^{N^{\text{SAM}}\times{W}\times{H}}}$, where $N^{\text{SAM}}$ is the number of potential target objects. We then multiply the vision mask $m_{i}\in{\{0,1\}^{W\times{H}}}$ of the ground truth action $i$ element-wise with each target object mask $m^{\text{SAM}}_{i}\in{\{0,1\}^{W\times{H}}}$ to obtain attention percentages $\mathbf{p}\in{\mathbb{R}}$ for all potential target objects. By re-ranking $\mathbf{p}$ and filtering out values below a predefined threshold, we derive the attention percentages on target objects. Finally, we manually label the target objects. Note that we preprocess the vision mask $m_{i}$ for attention percentage calculation by setting values below the average to zero. This preprocessing step filters out objects that received less attention but still show high attention percentages due to their large size, such as the background sky. The hyper-parameters used in SAM automatic mask generator: $\text{points\_per\_side}=64,\text{pred\_iou\_thresh}=0.7,\text{stability\_score\_thresh}=0.8$. The threshold we set on attention percentage calculation is $0.05$.